# On geometric spectral functionals

### v1

@@ 3 General results on spectral functionals
If $B=b_{a}\gamma^{a}$, where $b_{a}$ are endomorphisms commuting with the Clifford module (in particular, $B$ can be an image of a one-form in the Clifford algebra), then for every endomorphism $E$, $\mathscr{W}\!\mathrm{res}(EDD^{-2m})=0$, so that $D$ is spectrally closed (as defined in [7, Definition 5.5]).
Proof 3.2.
To demonstrate this, we compute the density depending on $B$ using Proposition 2.5:
$\mathrm{Tr}\left[E(2e_{b}\gamma^{b}-\gamma^{a}\{\gamma^{a},e_{c}\gamma^{c}\})\right]=\mathrm{Tr}\left(E\left(2e_{b}\gamma^{b}-e_{c}\gamma^{a}2\delta^{ac}\right)\right)=0.$ (3.1)
@@
If $E=e_{a}\gamma^{a}$, where $e_{a}$ are endomorphisms commuting with the Clifford module (in particular, $E$ can be an image of a one-form in the Clifford algebra), then the functional $\mathscr{W}\!\mathrm{res}(EDD^{-2m})$ vanishes identically for every $B$.
Proof 3.4.
@@
$\mathrm{Tr}\left[e_{b}\gamma^{b}(2B_{0}-\gamma^{a}\{\gamma^{a},B_{0}\})\right]$ $=\mathrm{Tr}\bigl{[}e_{b}\left(2\gamma^{b}B_{0}-\gamma^{b}\gamma^{a}\gamma^{a}B_{0}-\gamma^{b}\gamma^{a}B_{0}\gamma^{a}\right)\bigr{]}$ (3.2)
$=\mathrm{Tr}\bigl{[}e_{b}\left(2\gamma^{b}B_{0}-\{\gamma^{b},\gamma^{a}\}\gamma^{a}B_{0}\right)\bigr{]}=0.$
Next, we study the dependence of the spectral functionals on the perturbations $B$ of the Dirac operator $D_{0}$. We have,
@@
$\mathscr{g}_{D}(u,w)=\mathscr{g}_{D_{0}}(u,w)=\mathrm{dim}(V)\,\nu_{n-1}\int_{M}\mathrm{dvol}_{g}\,g(u,v).$ (3.3)
@@
For $D=D_{0}+B$ satisfying the assumptions of this paper, the Einstein functional density in normal coordinates reads:
$\mathscr{G}_{D}(u,w)(x)=$ $\mathscr{G}_{D_{0}}(u,w)(x)$ (3.4)
$+$ $\frac{\nu_{n-1}}{2}\mathrm{Tr}\Big{\{}iu_{a}w_{bc}[\gamma^{a},\gamma^{b}]\{\gamma^{c},B_{0}\}+\frac{i}{2}u_{a}w_{b}[\gamma^{a},\gamma^{b}]\{\gamma^{c},B_{c}\}$
$\hskip 20.0pt+u_{a}w_{b}\left[(\delta^{ab}B_{0}-\gamma^{a}\{\gamma^{b},B_{0}\})(\{\gamma^{c},B_{0}\}\gamma^{c}-2B_{0})\right]\Big{\}}.$
Proof 3.7.
The proof is identical to the first part in [3, Proposition II.1].
@@
Proof 3.9.
For $B^{\prime}=B+A_{a}\gamma^{a}$, we have $\{\gamma^{a},B^{\prime}-B\}=2A_{a}$, therefore $\{\gamma^{c},B_{0}^{\prime}-B_{0}\}\gamma^{c}-2(B_{0}^{\prime}-B_{0})=0$ and $\mathrm{Tr}[\gamma^{a},\gamma^{b}]\{\gamma^{c},B^{\prime}-B\}=0$. The last identity ensures that the first two linear terms in $B^{\prime}$ remain as they were for $B$. To complete the proof, it remains to show that the quadratic terms are the same for $B$ and $B^{\prime}=B+A_{a}\gamma^{a}$. Indeed, $\{\gamma^{c},B_{0}^{\prime}\}\gamma^{c}-2B_{0}^{\prime}=\{\gamma^{c},B_{0}\}\gamma^{c}-2B_{0}$, while
$(\delta^{ab}B_{0}^{\prime}-\gamma^{a}\{\gamma^{b},B_{0}^{\prime}\})-(\delta^{ab}B_{0}-\gamma^{a}\{\gamma^{b},B_{0}\})=\gamma^{c}(\delta^{ab}A_{c}-2\delta^{ac}A_{b}).$
The claim follows from the proof of Proposition 3.3.
@@
$\mathscr{T}_{D}(u,v,w)$ $=\nu_{n-1}\int_{M}\mathrm{dvol}_{g}\,u_{a}v_{b}w_{c}\mathrm{Tr}\left[(\gamma^{a}\gamma^{c}\gamma^{b}-\gamma^{b}\gamma^{c}\gamma^{a})B_{0}\right]$ (3.5)
$=\frac{\nu_{n-1}}{2}\int_{M}\mathrm{dvol}_{g}\,u_{a}v_{b}w_{c}\mathrm{Tr}\left([\gamma^{b},\gamma^{a}]\{\gamma^{c},B_{0}\}\right).$
@@
$\mathscr{T}_{D}(u,v,w)$ $=\frac{\nu_{n-1}}{2}\int_{M}\mathrm{dvol}_{g}\,u_{a}v_{b}w_{c}\mathrm{Tr}\left[\gamma^{a}\gamma^{b}\gamma^{c}(2B_{0}-\gamma^{d}\{\gamma^{d},B_{0}\})\right]$ (3.6)
$=\frac{\nu_{n-1}}{2}\int_{M}\mathrm{dvol}_{g}\,u_{a}v_{b}w_{c}\mathrm{Tr}\left[(2\gamma^{a}\gamma^{b}\gamma^{c}-\gamma^{a}\gamma^{b}\gamma^{c}\gamma^{d}\gamma^{d}-\gamma^{d}\gamma^{a}\gamma^{b}\gamma^{c}\gamma^{d})B_{0}\right]$
$=\frac{\nu_{n-1}}{2}\int_{M}\mathrm{dvol}_{g}\,u_{a}v_{b}w_{c}\mathrm{Tr}\left[(2\gamma^{a}\gamma^{b}\gamma^{c}-2\delta^{ad}\gamma^{b}\gamma^{c}\gamma^{d}+2\delta^{bd}\gamma^{a}\gamma^{c}\gamma^{d}-2\delta^{cd}\gamma^{a}\gamma^{b}\gamma^{c})B_{0}\right]$
$=\frac{\nu_{n-1}}{2}\int_{M}\mathrm{dvol}_{g}\,u_{a}v_{b}w_{c}\mathrm{Tr}\left[2(\gamma^{a}\gamma^{c}\gamma^{b}-\gamma^{b}\gamma^{c}\gamma^{a})B_{0}\right].$
The combination $\gamma^{a}\gamma^{c}\gamma^{b}-\gamma^{b}\gamma^{c}\gamma^{a}$ is totally antisymmetric in $a,b,c$, which is clear for the pair $a,b$. For other pairs, e.g., $a,c$, we can write:
$\gamma^{c}\gamma^{a}\gamma^{b}-\gamma^{b}\gamma^{a}\gamma^{c}=(2\delta^{ac}\gamma^{b}-\gamma^{a}\gamma^{c}\gamma^{b})-(2\gamma^{b}\delta^{ac}-\gamma^{b}\gamma^{c}\gamma^{a})=-(\gamma^{a}\gamma^{c}\gamma^{b}-\gamma^{b}\gamma^{c}\gamma^{a}).$ (3.7)
The last form (with $[\gamma^{b},\gamma^{a}]$) follows from splitting
$\gamma^{a}\gamma^{c}\gamma^{b}=\frac{1}{2}(2\delta^{ac}\gamma^{b}-\gamma^{c}\gamma^{a}\gamma^{b}+2\delta^{bc}\gamma^{a}-\gamma^{a}\gamma^{b}\gamma^{c}),$ (3.8)
and similarly for $\gamma^{b}\gamma^{c}\gamma^{a}$, then using the trace property.
@@
Finally, we study the impact of $B$ on the scalar curvature functional. Recall that the dependence of this functional on torsion has been extensively studied in some special cases; however, no general formula applicable, for instance, to the Hodge-Dirac operator has been obtained so far.
@@
$\begin{split}\mathscr{R}_{D}(f)=\mathscr{R}_{D_{0}}(f)+\frac{(n-2)\nu_{n-1}}{24}\int_{M}\mathrm{dvol}_{g}f\,\mathrm{Tr}\Bigl{(}-12B_{0}^{2}+6\gamma^{a}\{\gamma^{a},B_{0}\}B_{0}\Bigr{)}.\end{split}$ (3.9)
Proof 3.14.
To compute the spectral functional, we start by applying Lemma 2.2 to an endomorphism $E=f$, i.e., a function on $M$,
$\mathscr{R}_{D}(f)=\frac{(n-2)\nu_{n-1}}{24}\int_{M}\mathrm{dvol}_{g}f\,\mathrm{Tr}\Bigl{(}-12Q+6P_{aa}-2R-3S_{a}S_{a}\Bigr{)}.$ (3.10)
@@
$\begin{split}\mathrm{Tr}\Bigl{(}-12Q+6P_{aa}-2R-3S_{a}S_{a}\Bigr{)}\;=\;&\mathrm{Tr}\Bigl{(}-12Q^{0}+6P^{0}_{aa}\\ &-2R-12B_{0}^{2}-12i\gamma^{a}B_{a}+6i\{\gamma^{a},B_{a}\}+3\{\gamma^{a},B_{0}\}\{\gamma^{a},B_{0}\}\Bigr{)}\\ &=\mathrm{Tr}\Bigl{(}-12Q^{0}+6P^{0}_{aa}-2R-12B_{0}^{2}+3\{\gamma^{a},B_{0}\}\{\gamma^{a},B_{0}\}\Bigr{)},\end{split}$ (3.11)
@@
$\begin{split}\mathrm{Tr}(\{\gamma^{a},B_{0}\}\{\gamma^{a},B_{0}\})&=\mathrm{Tr}(\{\gamma^{a},B_{0}\}\gamma^{a}B_{0}+\{\gamma^{a},B_{0}\}B_{0}\gamma^{a})=\mathrm{Tr}(2\gamma^{a}\{\gamma^{a},B_{0}\}B_{0}),\end{split}$ (3.12)
which follows from the fact that $\{\gamma^{a},B_{0}\}\gamma^{a}=\gamma^{a}\{\gamma^{a},B_{0}\}$ and the cyclicity of the trace. Note that the final result does not depend on the derivatives of $B$. The part of the expression that does not depend on $B$ yields the value of the scalar curvature functional for $L=D_{0}^{2}$.
@@
Proof 3.16.
Recall that for $D^{2}_{0}$, we have (using the notation of Definition 2.4)
$P^{0}_{ab}=i\{\gamma^{a},g_{b}\}+2\gamma^{c}f^{a}_{bc},\qquad S_{a}^{0}=0,\qquad Q^{0}=i\gamma^{a}g_{a},$ (3.13)
@@
$\mathrm{Tr}\Bigl{(}-12Q^{0}+6P^{0}_{aa}-2R-3S^{0}_{a}S^{0}_{a}\Bigr{)}$ $=\mathrm{Tr}\Bigl{(}-12i\gamma^{a}g_{a}+6i\{\gamma^{a},g_{a}\}+12\gamma^{c}f^{a}_{ac}-2R\Bigr{)}$
$=\mathrm{Tr}\Bigl{(}6\{\gamma^{c},f^{a}_{ac}\}-2R\Bigr{)},$
@@
($\left(\{\gamma^{a},f^{b}_{cd}\}+\frac{1}{3}R_{acbd}\right)\xi_{a}\xi_{b}x^{c}x^{d}=0.$ (3.14)
This comes from the requirement that $D_{0}^{2}$ has the same expansion of the principal symbol as the Laplace operator. Substituting this into the previous equation yields the result.
@@ 3.1 The spin Dirac operator
As this case was the subject of a separate study [3] we start by briefly summarizing the main results concerning spectral functionals discussed therein, and slightly extend some of them.
@@
$D_{0}=i\gamma^{i}\nabla_{e_{i}}^{(s)}=i\gamma^{i}e_{i}-\frac{1}{4}\omega_{ijk}\gamma^{i}\gamma^{j}\gamma^{k},$ (3.15)
with the spin connection $\omega_{ijk}=\frac{1}{2}(c_{ijk}+c_{kij}+c_{kji})$ defined by structure constants $[e_{i},e_{j}]=c_{ijk}e_{k}$. Since in the normal coordinates
$c_{pqr}=\frac{1}{2}R_{pqnr}x^{n}+o({\bf x}),\qquad\omega_{ijk}=-\frac{1}{2}R_{nijk}x^{n}+o({\bf x}),\qquad\omega_{ijk}\gamma^{i}\gamma^{j}\gamma^{k}=\mathrm{Ric}_{ab}\gamma^{b}x^{a},$ (3.16)
@@
$D_{0}=i\gamma^{a}\Bigl{(}\partial_{a}-\frac{1}{6}R_{abcd}x^{b}x^{c}\partial_{d}-\frac{1}{4}\mathrm{Ric}_{ab}x^{b}\Bigr{)},$ (3.17)
@@
$D_{0}^{2}=$ $-\Bigl{(}\delta_{ab}+\frac{1}{3}R_{acbd}x^{c}x^{d}+o({\bf x}^{2})\Bigr{)}\partial_{a}\partial_{b}+\Biggl{[}\frac{1}{4}R_{kjba}\gamma^{j}\gamma^{k}+\frac{2}{3}\mathrm{Ric}_{ab}\Biggr{]}x^{b}\partial_{a}+\frac{1}{4}R.$ (3.18)
@@
On a spin manifold $M$ with the spinor bundle $S$ and $D_{0}$ the Dirac operator (3.17), if the perturbation $B$ is of the form,
$B=-\frac{i}{8}{}^{A}T_{ijk}\gamma^{i}\gamma^{j}\gamma^{k},$ (3.19)
with a totally antisymmetric tensor ${}^{A}T_{ijk}$, to which we refer as a torsion, we have that the density of the Einstein functional in normal coordinates is:
$\mathscr{G}_{D}(u,w)(x)-$ $\mathscr{G}_{D_{0}}(u,w)(x)=3\cdot 2^{n-1}\nu_{n-1}\Bigl{[}-u_{a}w_{bc}{}^{A}T^{0}_{abc}$ (3.20)
$\qquad\qquad+\frac{1}{8}u_{a}w_{b}\Bigl{(}\delta^{ab}{}^{A}T^{0}_{ijk}{}^{A}T^{0}_{ijk}-4{}^{A}T^{c}_{abc}-6{}^{A}T^{0}_{ajk}{}^{A}T^{0}_{bjk}\Bigr{)}\Bigr{]}.$
Proof 3.18.
@@
$\begin{split}&\{\gamma^{a},B\}=-\frac{3i}{4}{}^{A}T_{ajk}\gamma^{j}\gamma^{k},\qquad\gamma^{a}\{\gamma^{a},B\}=\{\gamma^{a},B\}\gamma^{a}=6B,\\ &\mathrm{Tr}(\gamma^{a}B)=0,\qquad\mathrm{Tr}(\gamma^{a}\gamma^{b}\{\gamma^{c},B\})=2\,\mathrm{Tr}(\gamma^{a}\gamma^{b}\gamma^{c}B)=-2^{m}\frac{3i}{2}{}^{A}T_{cba}.\end{split}$ (3.21)
Expanding ${}^{A}T_{ijk}={}^{A}T^{0}_{ijk}+{}^{A}T^{c}_{ijk}x^{c}+o({\bf x})$, and using Proposition 3.6, we end up with the form of the Einstein functional in the presence of torsion.
Note that the above functional constitutes a nontrivial contribution to the Einstein functional arising from the antisymmetric torsion and that the resulting density is not tensorial (the density of the functional is not $C^{\infty}(M)$–bilinear), as it depends on the derivatives of the forms. Furthermore, we note that the part of the functional that is non-tensorial remains symmetric in $u,w$ up to a total derivative,
$-u_{a}w_{bc}{}^{A}T^{0}_{abc}-\frac{1}{2}u_{a}w_{b}{}^{A}T^{c}_{abc}=-\partial_{c}\bigl{(}u_{a}w_{b}{}^{A}T_{abc}\bigr{)}-w_{a}u_{bc}{}^{A}T^{0}_{abc}-\frac{1}{2}w_{a}u_{b}{}^{A}T^{c}_{abc}.$ (3.22)

### v2

@@ 3 General results on spectral functionals
If $B=b_{a}\gamma^{a}$, where $b_{a}$ are endomorphisms commuting with the Clifford module ((in particular, $B$ can be an image of a one-form in the Clifford algebra)), then for every endomorphism $E$, $\mathscr{W}\mathrm{res}\bigl(EDD^{-2m}\bigr)=0$, so that $D$ is spectrally closed, that is for any zero-order operator $O$ the Wodzicki residuum of $OD|D|^{-2m}$ vanishes ((cf. [7, Definition 5.5]).
Proof.
To demonstrate this, we compute the density depending on $B$ using Proposition 2.5
$\mathrm{Tr}\bigl[E\bigl(2b_{b}\gamma^{b}-\gamma^{a}\{\gamma^{a},b_{c}\gamma^{c}\}\bigr)\bigr]=\mathrm{Tr}\bigl(E\bigl(2b_{b}\gamma^{b}-b_{c}\gamma^{a}2\delta^{ac}\bigr)\bigr)=0.$
@@
If $E=e_{a}\gamma^{a}$, where $e_{a}$ are endomorphisms commuting with the Clifford module ((in particular, $E$ can be an image of a one-form in the Clifford algebra)), then the functional $\mathscr{W}\mathrm{res}\bigl(EDD^{-2m}\bigr)$ vanishes identically for every $B$.
Proof.
@@
$\mathrm{Tr}\bigl[e_{b}\gamma^{b}(2B_{0}-\gamma^{a}\{\gamma^{a},B_{0}\})\bigr]$ $=\mathrm{Tr}\bigl[e_{b}\bigl(2\gamma^{b}B_{0}-\gamma^{b}\gamma^{a}\gamma^{a}B_{0}-\gamma^{b}\gamma^{a}B_{0}\gamma^{a}\bigr)\bigr]$
$=\mathrm{Tr}\bigl[e_{b}\bigl(2\gamma^{b}B_{0}-\bigl\{\gamma^{b},\gamma^{a}\bigr\}\gamma^{a}B_{0}\bigr)\bigr]=0.$
We remark that in Proposition 3.1, we assume the form of the endomorphism $B$ and compute the functional $\mathscr{W}\mathrm{res}\bigl(EDD^{-2m}\bigr)$ for arbitrary endomorphism $E$, while in Proposition 3.2, the form of $E$ is assumed and the result is stated for arbitrary $B$.
Next, we study the dependence of the spectral functionals on the perturbations $B$ of the Dirac operator $D_{0}$. We have the following.
@@
$\mathscr{g}_{D}(u,w)=\mathscr{g}_{D_{0}}(u,w)=\mathrm{dim}(V)\nu_{n-1}\int_{M}\mathrm{dvol}_{g}g(u,v).$
@@
For $D=D_{0}+B$ satisfying the assumptions of this paper ((cf. Definition 2.4), the Einstein functional density in normal coordinates reads
$\mathscr{G}_{D}(u,w)(x)={}$ $\mathscr{G}_{D_{0}}(u,w)(x)+\frac{\nu_{n-1}}{2}\mathrm{Tr}\biggl\{iu_{a}w_{bc}\big[\gamma^{a},\gamma^{b}\big]\{\gamma^{c},B_{0}\}+\frac{\rm i}{2}u_{a}w_{b}\bigl[\gamma^{a},\gamma^{b}\bigr]\{\gamma^{c},B_{c}\}$
$+u_{a}w_{b}\bigl[\bigl(\delta^{ab}B_{0}-\gamma^{a}\bigl\{\gamma^{b},B_{0}\bigr\}\bigr)(\{\gamma^{c},B_{0}\}\gamma^{c}-2B_{0})\bigr]\biggr\}.$
Proof.
The proof is identical to the first part in [2, Proposition II.1]. ∎
@@
Proof.
For $B^{\prime}=B+A_{a}\gamma^{a}$, we have $\{\gamma^{c},B^{\prime}-B\}=2A_{a}\delta^{ac}$, therefore $\{\gamma^{c},B_{0}^{\prime}-B_{0}\}\gamma^{c}-2(B_{0}^{\prime}-B_{0})=0$ and $\mathrm{Tr}\bigl[\gamma^{a},\gamma^{b}\bigr]\{\gamma^{c},B^{\prime}-B\}=0$. The last identity ensures that the first two linear terms in $B^{\prime}$ remain as they were for $B$. To complete the proof, it remains to show that the quadratic terms are the same for $B$ and $B^{\prime}=B+A_{a}\gamma^{a}$. Indeed, $\{\gamma^{c},B_{0}^{\prime}\}\gamma^{c}-2B_{0}^{\prime}=\{\gamma^{c},B_{0}\}\gamma^{c}-2B_{0}$, while
$\bigl(\delta^{ab}B_{0}^{\prime}-\gamma^{a}\bigl\{\gamma^{b},B_{0}^{\prime}\bigr\}\bigr)-\bigl(\delta^{ab}B_{0}-\gamma^{a}\bigl\{\gamma^{b},B_{0}\bigr\}\bigr)=\gamma^{c}\bigl(\delta^{ab}A_{c}-2\delta^{ac}A_{b}\bigr).$
The claim follows from the proof of Proposition 3.2. ∎
@@
$\mathscr{T}_{D}(u,v,w)$ $=\nu_{n-1}\int_{M}\mathrm{dvol}_{g}u_{a}v_{b}w_{c}\mathrm{Tr}\bigl[\bigl(\gamma^{a}\gamma^{c}\gamma^{b}-\gamma^{b}\gamma^{c}\gamma^{a}\bigr)B_{0}\bigr]$
$=\frac{\nu_{n-1}}{2}\int_{M}\mathrm{dvol}_{g}u_{a}v_{b}w_{c}\mathrm{Tr}\bigl(\bigl[\gamma^{b},\gamma^{a}\bigr]\{\gamma^{c},B_{0}\}\bigr).$
@@
$\mathscr{T}_{D}(u,v,w)={}$ $\frac{\nu_{n-1}}{2}\int_{M}\mathrm{dvol}_{g}u_{a}v_{b}w_{c}\mathrm{Tr}\bigl[\gamma^{a}\gamma^{b}\gamma^{c}(2B_{0}-\gamma^{d}\bigl\{\gamma^{d},B_{0}\bigr\})\bigr]$
$={}$ $\frac{\nu_{n-1}}{2}\int_{M}\mathrm{dvol}_{g}u_{a}v_{b}w_{c}\mathrm{Tr}\bigl[\bigl(2\gamma^{a}\gamma^{b}\gamma^{c}-\gamma^{a}\gamma^{b}\gamma^{c}\gamma^{d}\gamma^{d}-\gamma^{d}\gamma^{a}\gamma^{b}\gamma^{c}\gamma^{d}\bigr)B_{0}\bigr]$
$={}$ $\frac{\nu_{n-1}}{2}\int_{M}\mathrm{dvol}_{g}u_{a}v_{b}w_{c}\mathrm{Tr}$
$\times\bigl[\bigl(2\gamma^{a}\gamma^{b}\gamma^{c}-2\delta^{ad}\gamma^{b}\gamma^{c}\gamma^{d}+2\delta^{bd}\gamma^{a}\gamma^{c}\gamma^{d}-2\delta^{cd}\gamma^{a}\gamma^{b}\gamma^{c}\bigr)B_{0}\bigr]$
$={}$ $\frac{\nu_{n-1}}{2}\int_{M}\mathrm{dvol}_{g}u_{a}v_{b}w_{c}\mathrm{Tr}\bigl[2\bigl(\gamma^{a}\gamma^{c}\gamma^{b}-\gamma^{b}\gamma^{c}\gamma^{a}\bigr)B_{0}\bigr].$
The combination $\gamma^{a}\gamma^{c}\gamma^{b}-\gamma^{b}\gamma^{c}\gamma^{a}$ is totally antisymmetric in $a$, $b$, $c$, which is clear for the pair aabb. For other pairs, e.g., $a$, $c$, we can write
$\gamma^{c}\gamma^{a}\gamma^{b}-\gamma^{b}\gamma^{a}\gamma^{c}=\bigl(2\delta^{ac}\gamma^{b}-\gamma^{a}\gamma^{c}\gamma^{b}\bigr)-\bigl(2\gamma^{b}\delta^{ac}-\gamma^{b}\gamma^{c}\gamma^{a}\bigr)=-\bigl(\gamma^{a}\gamma^{c}\gamma^{b}-\gamma^{b}\gamma^{c}\gamma^{a}\bigr).$
The last form (with $\bigl[\gamma^{b},\gamma^{a}\bigr]$) follows from splitting
$\gamma^{a}\gamma^{c}\gamma^{b}=\frac{1}{2}\bigl(2\delta^{ac}\gamma^{b}-\gamma^{c}\gamma^{a}\gamma^{b}+2\delta^{bc}\gamma^{a}-\gamma^{a}\gamma^{b}\gamma^{c}\bigr),$
and similarly for $\gamma^{b}\gamma^{c}\gamma^{a}$, then using the trace property. ∎
@@
Finally, we study the impact of $B$ on the scalar curvature functional. Recall that the dependence of this functional on torsion has been extensively studied in some special cases, where torsion enters through Dirac operators and contributes additional terms to heat-kernel coefficients and the spectral action [14, 17]; however, no general formula applicable, for instance, to the Hodge–Dirac operator has been obtained so far.
@@
$\mathscr{R}_{D}(f)=\mathscr{R}_{D_{0}}(f)+\frac{(n-2)\nu_{n-1}}{24}\int_{M}\mathrm{dvol}_{g}f\mathrm{Tr}\bigl(-12B_{0}^{2}+6\gamma^{a}\{\gamma^{a},B_{0}\}B_{0}\bigr).$
Proof.
To compute the spectral functional, we start by applying Corollary 2.2 to an endomorphism $E=f$, i.e., a function on $M$,
$\mathscr{R}_{D}(f)=\frac{(n-2)\nu_{n-1}}{24}\int_{M}\mathrm{dvol}_{g}f\mathrm{Tr}(-12Q+6P_{aa}-2R-3S_{a}S_{a}).$
@@
$\mathrm{Tr}(-12Q+6P_{aa}-2R-3S_{a}S_{a})$
$\qquad=\mathrm{Tr}\bigl(-12Q^{0}+6P^{0}_{aa}-2R-12B_{0}^{2}-12i\gamma^{a}B_{a}+6i\{\gamma^{a},B_{a}\}+3\{\gamma^{a},B_{0}\}\{\gamma^{a},B_{0}\}\bigr)$
$\qquad=\mathrm{Tr}\bigl(-12Q^{0}+6P^{0}_{aa}-2R-12B_{0}^{2}+3\{\gamma^{a},B_{0}\}\{\gamma^{a},B_{0}\}\bigr),$
@@
$\mathrm{Tr}(\{\gamma^{a},B_{0}\}\{\gamma^{a},B_{0}\})=\mathrm{Tr}(\{\gamma^{a},B_{0}\}\gamma^{a}B_{0}+\{\gamma^{a},B_{0}\}B_{0}\gamma^{a})=\mathrm{Tr}(2\gamma^{a}\{\gamma^{a},B_{0}\}B_{0}),$
which follows from the fact that $\{\gamma^{a},B_{0}\}\gamma^{a}=\gamma^{a}\{\gamma^{a},B_{0}\}$ and the cyclicity of the trace. Note that the final result does not depend on the derivatives of $B$. The part of the expression that does not depend on $B$ yields the value of the scalar curvature functional for $L=D_{0}^{2}$. ∎
@@
Proof.
Recall that for $D^{2}_{0}$, we have (cf. Definition 2.4)
$P^{0}_{ab}={\rm i}\{\gamma^{a},g_{b}\}+2\gamma^{c}f^{a}_{bc},\qquad S_{a}^{0}=0,\qquad Q^{0}={\rm i}\gamma^{a}g_{a},$
@@
$\mathrm{Tr}\bigl(-12Q^{0}+6P^{0}_{aa}-2R-3S^{0}_{a}S^{0}_{a}\bigr)$ $=\mathrm{Tr}(-12{\rm i}\gamma^{a}g_{a}+6{\rm i}\{\gamma^{a},g_{a}\}+12\gamma^{c}f^{a}_{ac}-2R)$
$=\mathrm{Tr}(6\{\gamma^{c},f^{a}_{ac}\}-2R),$
@@
$\left(\bigl\{\gamma^{a},f^{b}_{cd}\bigr\}+\frac{1}{3}R_{acbd}\right)\xi_{a}\xi_{b}x^{c}x^{d}=0.$
This comes from the requirement that $D_{0}^{2}$ has the same expansion of the principal symbol as the Laplace operator. Substituting this into the previous equation yields the result. ∎
@@ 3.1 The spin Dirac operator
As this case was the subject of a separate study [2], we start by briefly summarizing the main results concerning spectral functionals discussed therein, and slightly extend some of them.
@@
$D_{0}={\rm i}\gamma^{i}\nabla_{e_{i}}^{(s)}={\rm i}\gamma^{i}e_{i}-\frac{1}{4}\omega_{ijk}\gamma^{i}\gamma^{j}\gamma^{k},$
with the spin connection $\omega_{ijk}=\frac{1}{2}(c_{ijk}+c_{kij}+c_{kji})$ defined by structure constants $e_{i},e_{j}]=c_{ijk}e_{k}$. Since in the normal coordinates
$c_{pqr}=\frac{1}{2}R_{pqnr}x^{n}+o({\bf x}),\qquad\omega_{ijk}=-\frac{1}{2}R_{nijk}x^{n}+o({\bf x}),\qquad\omega_{ijk}\gamma^{i}\gamma^{j}\gamma^{k}=\mathrm{Ric}_{ab}\gamma^{b}x^{a},$
@@
$D_{0}={\rm i}\gamma^{a}\left(\partial_{a}-\frac{1}{6}R_{abcd}x^{b}x^{c}\partial_{d}-\frac{1}{4}\mathrm{Ric}_{ab}x^{b}\right),$ (3.1)
@@
$D_{0}^{2}=$ $-\left(\delta_{ab}+\frac{1}{3}R_{acbd}x^{c}x^{d}+o\bigl({\bf x}^{2}\bigr)\right)\partial_{a}\partial_{b}+\left[\frac{1}{4}R_{kjba}\gamma^{j}\gamma^{k}+\frac{2}{3}\mathrm{Ric}_{ab}\right]x^{b}\partial_{a}+\frac{1}{4}R.$
@@
On a spin manifold $M$ with the spinor bundle $S$ and $D_{0}$ the Dirac operator (3.1), if the perturbation $B$ is of the form
$B=-\frac{\rm i}{8}{}^{A}T_{ijk}\gamma^{i}\gamma^{j}\gamma^{k},$ (3.2)
with a totally antisymmetric tensor $^{A}T_{ijk}$, to which we refer as a torsion, we have that the density of the Einstein functional in normal coordinates is
$\mathscr{G}_{D}(u,w)(x)-\mathscr{G}_{D_{0}}(u,w)(x)={}$ $3\cdot 2^{n-1}\nu_{n-1}\biggl[-u_{a}w_{bc}{}^{A}T^{0}_{abc}$
$+\frac{1}{8}u_{a}w_{b}\bigl(\delta^{ab}{}^{A}T^{0}_{ijk}{}^{A}T^{0}_{ijk}-4{}^{A}T^{c}_{abc}-6{}^{A}T^{0}_{ajk}{}^{A}T^{0}_{bjk}\bigr)\biggr],$
we use the expansion $^{A}T_{ijk}={}^{A}T^{0}_{ijk}+{}^{A}T^{c}_{ijk}x^{c}+o({\bf x})$.
Proof.
@@
$\{\gamma^{a},B\}=-\frac{3{\rm i}}{4}{}^{A}T_{ajk}\gamma^{j}\gamma^{k},\qquad\gamma^{a}\{\gamma^{a},B\}=\{\gamma^{a},B\}\gamma^{a}=6B,$
$\mathrm{Tr}(\gamma^{a}B)=0,\qquad\mathrm{Tr}\bigl(\gamma^{a}\gamma^{b}\{\gamma^{c},B\}\bigr)=2\mathrm{Tr}\bigl(\gamma^{a}\gamma^{b}\gamma^{c}B\bigr)=-2^{m}\frac{3{\rm i}}{2}{}^{A}T_{cba}.$ (3.3)
Using then Proposition 3.4, we end up with the form of the Einstein functional in the presence of torsion. ∎
Note that the above functional constitutes a nontrivial contribution to the Einstein functional arising from the antisymmetric torsion and that the resulting density is not tensorial (the density of the functional is not $C^{\infty}(M)$-bilinear), as it depends on the derivatives of the forms. Furthermore, we note that the part of the functional that is non-tensorial remains symmetric in uuww up to a total derivative,
$u_{a}w_{bc}{}^{A}T^{0}_{abc}-\frac{1}{2}u_{a}w_{b}{}^{A}T^{c}_{abc}=-\partial_{c}\bigl(u_{a}w_{b}{}^{A}T_{abc}\bigr)-w_{a}u_{bc}{}^{A}T^{0}_{abc}-\frac{1}{2}w_{a}u_{b}{}^{A}T^{c}_{abc}.$